OpenStudy (anonymous):

This is my question: Using the definition of the derivative at a point find f'(3) if f(x)=3x^3-2x^2+5. (Hint synthetic division may be involved)

OpenStudy (anonymous):

I'm not sure if this is right but: Find the derivative by using differentiation: f(x) = 3x^3-2x^2+5 (multiply the coefficient by the index and then subtract 1 from it) f'(x) = (3*3)x^(3-1) - (2*2)x^(2-1) + 5x^(1-1) f'(x) = 9x^2-4x So now we can find f'(3): f'(3) = 9*3^2-4*3 f'(3) = 81-12 f'(3) = 69 Therefore f'(3) = 69

OpenStudy (anonymous):

thanks that looks right!